Solve the systems of equations graphically and check x-y=5 y=-2x+4
Answer by ewatrrr(24785)

x-y=5 0r y = x-5
y=-2x+4
0r
y = x-5
P(0,-5) and P(2, -3) on this line: Plot and connect with the Line.
y=-2x+4
P(0, 4) and (2,0) on this line: Plot and connect with the Line.
P(3,-2) is the ordered pair that is the solution for this system of EQs
On may CHECK by substituting x= 3 and y = -2 into the EQs of these Lines